OpenStudy (anonymous):

-8x + 9x - 21 = -12 -16 Solve for x

OpenStudy (sleepyjess):

Welcome to OpenStudy!

OpenStudy (sleepyjess):

First we need to get - 21 to the right side

jagr2713 (jagr2713):

\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] @koolkat13 what do you think we do first?

OpenStudy (anonymous):

find out the -12 - 16 which is -28

OpenStudy (sleepyjess):

Yep, that is correct. How do you think we get -21 to the right side?

OpenStudy (anonymous):

idk

OpenStudy (sleepyjess):

What is the opposite of subtracting?

OpenStudy (anonymous):

adding

OpenStudy (sleepyjess):

Yes, so add 21 to both sides.

OpenStudy (anonymous):

ok so now we have -8x + -9x = -7

OpenStudy (sleepyjess):

Just 8x + 9x = -7 :) Combine 8x and 9x

OpenStudy (sleepyjess):

Oh, -8x + 9x

OpenStudy (anonymous):

17x my computer is freezing up....grr

OpenStudy (sleepyjess):

I corrected myself, it should be -8x+9x, or 9x-8x, either way

OpenStudy (anonymous):

ok then 1x?

OpenStudy (sleepyjess):

Yes! So x = -7

OpenStudy (anonymous):

oh ok thank you!...um do you have time for a few more?

OpenStudy (sleepyjess):

Sure, just close this question and open a new one :)

OpenStudy (sleepyjess):

You can also click "Best Response" to let someone know that they helped you.

OpenStudy (anonymous):

ok :)

OpenStudy (sleepyjess):

When you open a new question, you can also tag me by putting @sleepyjess in the comment box